A 4-foot tall girl casts a shadow that is 7.2 feet long. A nearby tree casts a shadow that is 25.56 feet long. How tall, in feet, is the tree?
Answer by stanbon(75887)

A 4-foot tall girl casts a shadow that is 7.2 feet long. A nearby tree casts a shadow that is 25.56 feet long. How tall, in feet, is the tree?
-----------------------
Proportion:
tree/25.56 = 4/7.2
tree = 25.56(4/7.2) = 14.2 ft -----------------
Cheers,
Stan H.--------------------